OpenStudy (pphalke):

For exercises 1-17, be sure to label your answers. Use pi=3.14, and round your answers to the nearest tenth. #5 find the area and circumference with radius 12 meters

OpenStudy (rsmith6559):

c = 2 * r * pi a = pi * r^2

OpenStudy (pphalke):

c=452.2 and a=75.4 does that sound right? I plugged numbers in

OpenStudy (rsmith6559):

I think you have them backwards. An area of 452.2 sounds reasonable, and so does a circumference of 75.4.